6x-6y=-48
-10x+4y=92

Answer by ewatrrr(24785)

 
Hi
x - y = -8 times 4 and then ADD to eliminate y
-10x + 4y = 92
-6x = 60
x = -10 and y = -2 -10-y=+-8